​DFB, Inc., expects earnings this year of $ 4.62 per​ share, and it plans to pay a $ 2.53 dividend to shareholders. DFB will retain $ 2.09 per share of its earnings to reinvest in new projects with an expected return of 15.5 % per year. Suppose DFB will maintain the same dividend payout​ rate, retention​ rate, and return on new investments in the future and will not change its number of outstanding shares.

a. What growth rate of earnings would you forecast for​ DFB?

b. If​ DFB's equity cost of capital is 12.9 % ​, what price would you estimate for DFB​ stock?

c. Suppose DFB instead paid a dividend of $ 3.53 per share this year and retained only $ 1.09 per share in earnings. That​ is, it chose to pay a higher dividend instead of reinvesting in as many new projects. If DFB maintains this higher payout rate in the​ future, what stock price would you estimate​ now? Should DFB follow this new​ policy?
Answer #1

(a) Current Earnings = E0 = $ 4.62, Current Dividend = D0 = $ 2.53 and Current Retained Earnings = $ 2.09

Retention Ratio = (2.09/4.62) = RR = 0.4524

Return on Reinvestment = r = 15.5 %

Growth Rate Generated = RR x r = g = 0.4524 x 15.5 = 7.012 %

(b) Equity Cost of Capital = ke = 12.9%

Current Stock Price = [D0 x (1+g)]/[ke-g] = [(2.53 x 1.07012) / (0.129 - 0.07012)] = $ 45.9809 ~ $ 45.98

(c) Current Dividend = D0 = $ 3.53 and Retained Earnings = $ 1.09

Retention Ratio = (1.09/4.62) = 0.23593

r = 15.5 %

Growth Generated = g = r x RR = 15.5 x 0.23593 = 3.65693 %

Current Stock Price = [D0 x (1+g) / (ke-g)] = [(3.53 x 1.0365693) / (0.129-0.0365693)] = $ 39.5874 ~ $ 39.59